# Sim-to-real causal transfer: A metric learning approach to causally-aware interaction representations

### v1

@@ 3.3 Diagnostic Dataset through Counterfactuals
Counterfactual Pairs.
To address the above caveats, we create a new diagnostic dataset through counterfactual simulations with ORCA (van den Berg et al., 2008). Specifically, we annotate the ground-truth causal effect by explicitly comparing the trajectories in paired scenes before and after agent removals, as defined in Eq. 2. Details of the collected data are summarized in Appendix B.
Fine-grained Category.
In addition to real-valued causal effects, we also seek to annotate the category of each agent. Specifically, from simulations, we extract the per-step causal relation between a neighboring agent $i$ and the ego agent. If agent $i$ is visible to the ego at time step $t$, it directly influences the ego, indicated by $\mathbbm{1}^{i}_{t}=1$; otherwise $\mathbbm{1}^{i}_{t}=0$. We then convert the causal effect over the whole sequence $\mathcal{E}_{i}$ into three agent categories:
@@ 4.1 Causally-Aware Regularization
Recall that the causal effect $\mathcal{E}_{i}$ of an agent $i$ is tied to the difference of the potential outcomes between the factual scene and the counterfactual one where the agent $i$ is removed. In this light, a causally-aware representation should also capture such relations, i.e., feature vectors of paired scenes are separated by a certain distance $d_{i}$ depending on $\mathcal{E}_{i}$. Motivated by this intuition, we measure the distance between a pair of counterfactual scenes in an embedding space through cosine distance,
$d_{i}=1-\textrm{sim}(\bm{p}_{\varnothing},\bm{p}_{i})=1-\bm{p}_{\varnothing}^{\top}\bm{p}_{i}/\lVert\bm{p}_{\varnothing}\rVert\lVert\bm{p}_{i}\rVert,$ (4)
where $\bm{p}_{i}=h(\bm{z}_{i})=h(f(\bm{x}_{i}))$ is a low-dimensional feature vector projected from the latent representation $\bm{z}_{i}$ through a non-linear head $h$. The desired representation is thus expected to preserve the following property,
@@
Causal Contrastive Regularization.
As discussed in Sec. 3.3, one simple form of causal annotations is binary labels, indicating whether or not a neighboring agent causally influences the behavior of the ego agent. Intuitively, the representations of paired counterfactual scenes with respect to non-causal agents should be quite different from those with respect to causal ones: the former should stay close to the factual scenes in the embedding space, whereas the latter should be rather far away. We formulate this intuition into a causal contrastive learning objective,
@@
where the positive (distant) example is sampled from counterfactual pairs with respect to causal agents, the negative (nearby) examples are sampled from counterfactual pairs with respect to non-causal agents, $\tau$ is a temperature hyperparameter controlling the difficulty of the contrastive task.
Causal Ranking Regularization.
One downside of causal contrastive learning described above is that it inherently ignores the detailed effect of causal agents. It tends to push the embeddings of counterfactuals equally far apart across all causal agents, regardless of the variation of causal effects, which violates the desired property stated in Eq. 5. To address this limitation, we further consider another variant of causal regularization using real-valued annotations to provide more dedicated supervision on the relative distance in the embedding space. Concretely, we first sort all agents in a scene based on their causal effect and then sample two agents with different causal effects for comparison. This allows us to formulate a ranking problem in a pairwise manner through a margin ranking loss,
@@ 4.2 Sim-to-real Causal Transfer
The causal regularization method described above relies upon the premise that annotations of causal effects are readily available. However, as elaborated in Sec. 3.3, procuring such causal annotations in real-world scenarios can be highly difficult. To bridge this gap, we extend our causal regularization approach to a sim-to-real transfer learning framework. Our key idea is that, despite discrepancies between simulation environments and the real world, some underlying causal mechanisms like collision avoidance and group coordination are likely stable across domains. As such, in the absence of real-world causal annotations, we can instead leverage the causal annotations derived from the simulation counterparts to jointly train the model on the prediction task $\mathcal{L}_{task}^{real}$ on the real-world data and the causal distance task on the simulation data $\mathcal{L}_{causal}^{syn}$,
$\mathcal{L}=\mathcal{L}_{\mathrm{task}}^{\mathrm{real}}+\alpha\mathcal{L}_{\mathrm{causal}}^{\mathrm{syn}},$ (8)
where $\alpha$ is a hyperparameter controlling the emphasis on the causal task. Despite its simplicity, we will show in Sec. 5.3 that the sim-to-real causal transfer framework effectively translates the causal knowledge absorbed from simulations to real-world benchmarks, e.g., the ETH-UCY dataset (Lerner et al., 2007; Pellegrini et al., 2010), even without any causal annotations from the latter.
@@ 5 Experiments
Throughout our experiments, the multi-agent forecasting task is defined as predicting the future trajectory of the ego agent for 12 time steps, given the history observations of all agents in a scene in the past 8 time steps. We evaluate forecasting models on three metrics:
  • Average Displacement Error (ADE): the average Euclidean distance between the predicted output and the ground truth trajectories – a metric widely used in prior literature to measure the prediction accuracy of a forecasting model.
@@
  • Average Causal Error (ACE): the average difference between the estimated causal effect and the ground truth causal effect – a metric inspired by Roelofs et al. (2022) to measure the causal awareness of a learned representation,
@@
Alongside the aggregated ACE, we also measure causal awareness on each category separately, i.e., ACE-NC, ACE-DC, ACE-IC for non-causal, direct causal, and indirect causal agents, respectively.
Table 1: Performance of modern representations on the created diagnostic dataset. The prediction errors and causal errors are generally substantial across all evaluated models. However, the errors associated with ACE-NC are rather marginal, suggesting that recent models are already partially robust to non-causal perturbations.
D-LSTM (Kothari et al., 2021) 0.329 0.677 0.027 0.532 0.614
S-LSTM (Alahi et al., 2016) 0.314 0.627 0.031 0.463 0.523
Trajectron++ Salzmann et al. (2020) 0.312 0.630 0.024 0.479 0.568
STGCNN (Mohamed et al., 2020) 0.307 0.564 0.049 0.330 0.354
AutoBots (Girgis et al., 2021) 0.255 0.497 0.045 0.595 0.616
@@ 5.1 Robustness of Recent Representations
Setup. We start our experiments with the evaluation of a collection of recent models on the diagnostic dataset described in Sec. 3.3, including S-LSTM (Alahi et al., 2016), D-LSTM (Kothari et al., 2021), Trajectron++ (Salzmann et al., 2020), STGCNN (Mohamed et al., 2020), and AutoBots (Girgis et al., 2021). Thanks to the fine-grained annotations from counterfactual simulations, we explicitly examine the errors of causal effect estimation for each agent category. We summarize the results in Tab. 1 and visualize the detailed estimation in Fig. 5.
Takeaway 1: recent representations are already partially robust to non-causal agent removal. As shown in Tab. 1, the values of ACE-NC are generally quite minimal compared to the results on other metrics. Across all evaluated models, the errors made upon non-causal agents are 10x smaller than that for causal agents (ACE-DC/IC). Corroborate with our analysis in Sec. 3.3, this result provides a counterargument to the robustness issue benchmarked in  Roelofs et al. (2022), suggesting the importance of studying the robustness under causal perturbations as opposed to non-causal ones.
@@ 5.2 Effectiveness of Causal Regularization
We next evaluate the efficacy of our causal regularization method on Autobots, the strongest baselines in Tab. 1. We take the data augmentation strategy proposed in (Roelofs et al., 2022) as a baseline, and compare different methods in two aspects: in-distribution causal awareness and out-of-distribution generalization.
(a) In-Distribution Evaluation
(b) Out-of-Distribution Evaluation
Figure 6: Quantitaive results of our causal regularization method in comparison to the Autobots baseline (Girgis et al., 2021) and the non-causal data augmentation (Roelofs et al., 2022) Models trained by our ranking-based method yield enhanced accuracy in estimating causal effects and more reliable predictions on out-of-distribution test sets. Results are averaged over five random seeds.
(a) ID
(b) OOD-Density
(c) OOD-Context
Figure 7: Qualitative results of our method on in-distribution (ID) and out-of-distribution (OOD) test sets. Models regularized by our ranking-based method demonstrate more robust understanding of agent interactions in scenarios with higher agent density and unseen scene context.
@@ 5.2.1 In-distribution Causal Awareness
Takeaway: our method boosts causal awareness thanks to fine-grained annotations. As shown in Fig. 5(a), both the contrastive and ranking variants of our causal regularization approach result in lower causal errors than the vanilla baseline. In particular, the causal ranking method demonstrates substantial advantages over the other counterparts across all quantities of causal effects, confirming the promise of incorporating fine-grained annotations in learning causally-aware representations.
@@ 5.2.2 Out-of-Distribution Generalization
Setup. We further examine the efficacy of our method by measuring the prediction accuracy on out-of-distribution test sets. We consider three common types of distribution shifts in the multi-agent setting: higher density, unseen context, and both combined. We summarize the results of different methods on the OOD test sets in Fig. 5(b) and visualize the difference of prediction output in Fig. 7.
Takeaway: causally-aware representations lead to stronger out-of-distribution generalization. As shown in Fig. 5(b), the prediction error from our causal ranking method is generally lower than that of the other methods on the OOD sets. In fact, the overall patterns between Fig. 5(a) and Fig. 5(b) are highly similar, indicating a substantial correlation between causal awareness and out-of-distribution robustness. This practical benefit is also visually evident in the qualitative visualization in Fig. 7.
@@ 5.3 Effectiveness of Causal Transfer
Setup. Finally, we evaluate the proposed causal transfer approach on the ETH-UCY dataset (Lerner et al., 2007; Pellegrini et al., 2010) paired with our diagnostic dataset. To assess the potential of the transfer method in challenging settings such as low-data regimes, we train Autobots (Girgis et al., 2021) on different fractions of the real-world data. The results are summarized in Fig. 8.
Takeaway: causal transfer improves performance in the real world despite domain gaps. As shown in Fig. 8, our causal transfer approach consistently improves the performance of the learned representation on the real-world benchmark. In particular, it enables the model to learn faster, e.g., even with half data in the real-world, our ranking-based method can still outperform other counterparts. In contrast, the vanilla sim-to-real technique (Liang et al., 2020) fails to yield effective transfer, especially in the low-data regimes. We conjecture that this is due to the disparity between simulation and the real world in terms of non-causal interaction styles, such as speed and curvature.
Figure 8: Results of causal transfer from ORCA simulations to the ETH-UCY dataset. We consider three settings where the model has access to the same simulation data but varying amounts of real-world data. Our ranking-based causal transfer results in lower prediction errors and higher learning efficiency than the other counterparts, i.e., AutoBots baseline (Girgis et al., 2021), the vanilla sim-to-real transfer (Liang et al., 2020) and the non-causal data augmentation (Roelofs et al., 2022). Details are summarized in Appendices B and A
Figure 8: Results of causal transfer from ORCA simulations to the ETH-UCY dataset. We consider three settings where the model has access to the same simulation data but varying amounts of real-world data. Our ranking-based causal transfer results in lower prediction errors and higher learning efficiency than the other counterparts, i.e., AutoBots baseline (Girgis et al., 2021), the vanilla sim-to-real transfer (Liang et al., 2020) and the non-causal data augmentation (Roelofs et al., 2022). Details are summarized in Appendices B and A
@@ 6 Conclusions
Summary. In this paper, we presented a thorough analysis and an effective approach for learning causally-aware representations of multi-agent interactions. We cast doubt on the notion of non-causal robustness in a recent benchmark (Roelofs et al., 2022) and showed that the main weaknesses of recent representations are not overestimations of the effect of non-causal agents but rather underestimation of the effects of indirect causal agents. To boost causal awareness of the learned representations, we introduced a regularization approach that encapsulates a contrastive and a ranking variant leveraging annotations of different granularity. We showed that our approach enables recent models to learn faster, generalize better, as well as transfer stronger to practical problems, even without real-world annotations.
Limitations. As one of the first steps towards causally-aware representations in the multi-agent context, our work is subject to two major limitations. On the technical front, while our proposed regularization method consistently boosts causal awareness in various settings, supervision alone is likely insufficient to fully solve causal reasoning in complex scenes, as evidenced by the causal errors in Fig. 5(a). Incorporating structural inductive biases might be a promising direction to address high-order reasoning challenges in the presence of indirect causal effects. On the empirical side, while we have demonstrated the potential of our approach in real-world settings in Sec. 5.3, most of our other experiments were conducted in controlled simulations for proof-of-concept. Scaling our findings to other practical benchmarks (e.g., WOMD (Ettinger et al., 2021) and INTERACTION (Zhan et al., 2019)) and contexts (e.g., navigation (Chen et al., 2019) and manipulation (Zhou et al., 2023)) can be another fruitful avenue for future research.
Acknowledgments
This work is supported by the Swiss National Science Foundation under the Grant 2OOO21-L92326.

### v2

@@ 3.3 Diagnostic Dataset through Counterfactuals
To address the identified issues, we have developed a new diagnostic dataset using the ORCA [65] simulation system, which enables the creation of counterfactual scenarios by removing sets of agents. ORCA models social interactions like collision avoidance and leader-follower behavior by taking initial positions, speeds, and final goals of agents. We constructed our dataset by initializing the positions and goals of $8\leq K+1\leq 18$ agents within a structured layout typical of an open area. Additionally, we generated two out-of-distribution (OOD) datasets, named Density OOD and Context OOD, to test model robustness. The Density OOD dataset varies the density of agents and adds more non-causal ones, while the Context OOD dataset alters the initial layout to resemble a narrow street with pedestrians moving from one end to the other. Both the datasets and the code used for their generation are available to support further research. For more detailed information on the data collected, see Appendix B.
Counterfactual Pairs. Leveraging the capability to simulate counterfactual scenarios, we annotate the ground-truth causal effects by comparing ego trajectories in paired scenes before and after agent removals, as defined in Eq. 2. Given the exponential growth of simulation scenarios with the number of agents, we adopt the approach of Causal Agents [56] and concentrate on evaluating causal effects at the level of individual agents, denoted as i.e. $|\mathcal{R}|=1$. For simplicity, moving forward, we will use $\mathcal{E}_{i}$ to refer to $\mathcal{E}_{\{i\}}$.
Fine-grained Category. In addition to real-valued causal effects, we also seek to annotate the category of each agent. From simulations, we extract the per-step causal relation between a neighboring agent $i$ and the ego agent. If agent $i$ is visible to the ego at time step $t$, it directly influences the ego, indicated by $\mathbbm{1}^{i}_{t}=1$; otherwise $\mathbbm{1}^{i}_{t}=0$. We then convert the causal effect over the whole sequence $\mathcal{E}_{i}$ into three agent categories:
@@ 4.1 Causally-Aware Regularization
Recall that the causal effect $\mathcal{E}_{i}$ of an agent $i$ is tied to the difference of the potential outcomes between the factual scene and the counterfactual one where the agent $i$ is removed. In this light, a causally-aware representation should also capture such relations, i.e., feature vectors of paired scenes are separated by a certain distance $d_{i}$ depending on $\mathcal{E}_{i}$.
We start by defining $\bm{p}_{\varnothing}$ as the low-dimensional feature vector of the factual scene where no agents are removed. This serves as our baseline scenario against which all counterfactual instances are compared. $\bm{p}_{\varnothing}$ is extracted from the latent representation $\bm{z}_{\varnothing}=f(\bm{x}_{\varnothing})$ projected through a non-linear head $h$. Similarly, $\bm{p}_{i}=h(\bm{z}_{i})=h(f(\bm{x}_{i}))$ the feature vector for the counterfactual scene $\bm{x}_{i}$ created by removing agent $i$. We then measure the distance between $\bm{x}_{\varnothing}$ and $\bm{x}_{i}$ as
$d_{i}=1-\textrm{sim}(\bm{p}_{\varnothing},\bm{p}_{i})=1-\frac{\bm{p}_{\varnothing}^{\top}\bm{p}_{i}}{\lVert\bm{p}_{\varnothing}\rVert\lVert\bm{p}_{i}\rVert},$ (4)
where $\textrm{sim}(\bm{p}_{\varnothing},\bm{p}_{i})$ denotes the cosine similarity between the feature vectors of the factual and counterfactual scenes, effectively measuring the impact of agent $i$’s removal on the embedding space. The desired representation is thus expected to preserve the following property,
@@
Causal Contrastive Regularization. As discussed in Sec. 3.3, one simple form of causal annotations is binary labels, indicating whether or not a neighboring agent causally influences the behavior of the ego agent. Intuitively, the embedding of a counterfactual scenario where a non-causal agent is removed should closely resemble the embedding of the factual scenario, reflecting minimal change from the ego agent’s perspective. In contrast, removing a causal agent should significantly alter the representation, indicating a higher impact. We formulate this intuition into a causal contrastive learning objective,
@@
where the positive (distant) example is sampled from counterfactuals with respect to causal agents, the negative (nearby) examples are sampled from counterfactual scenarios with respect to non-causal agents, and $\tau$ is a temperature hyperparameter controlling the difficulty of the contrastive task.
Causal Ranking Regularization. One downside of causal contrastive learning described above is that it inherently ignores the detailed effect of causal agents. It tends to push the embeddings of counterfactuals equally far apart across all causal agents, regardless of the variation of causal effects, which violates the desired property stated in Eq. 5. To address this limitation, we further consider another variant of causal regularization using real-valued causal annotations to provide more dedicated supervision on the relative distance in the embedding space. Concretely, we first sort all agents in a scene based on their causal effect and then sample two agents with different causal effects for comparison. This allows us to formulate a ranking problem in a pairwise manner through a margin ranking loss,
@@
D-LSTM [32] 0.329 0.677 0.027 0.532 0.614
S-LSTM [2] 0.314 0.627 0.031 0.463 0.523
Trajectron++ [61] 0.312 0.630 0.024 0.479 0.568
STGCNN [49] 0.307 0.564 0.049 0.330 0.354
Multi-Transmotion [18] 0.237 0.466 0.031 0.723 0.724
AutoBots [22] 0.255 0.497 0.045 0.595 0.616
Table 1: Performance of modern representations on the created diagnostic dataset. The prediction errors and causal errors are generally considerable across all evaluated models. However, the errors associated with ACE-NC are rather marginal, suggesting that recent models are already partially robust to non-causal perturbations of individual agents.
@@ 4.2 Sim-to-real Causal Transfer
The causal regularization method described above relies upon the premise that annotations of causal effects are readily available. However, as elaborated in Sec. 3.3, procuring such causal annotations in real-world scenarios can be difficult. To bridge this gap, we extend our causal regularization approach to a sim-to-real transfer learning framework. Our key idea is that, despite the gap between simulation environments and the real world, certain causal mechanisms (e.g., collision avoidance and group coordination) tend to remain stable. As such, in the absence of real-world causal annotations, we can use the annotations gathered from simulations as a viable alternative. This leads to cross-domain training on two tasks jointly: the main prediction task $\mathcal{L}_{pred}^{real}$ in real-world scenarios which usually is a reconstruction loss, and our auxiliary causal task $\mathcal{L}_{causal}^{syn}$ in the simulation counterpart,
$\mathcal{L}=\mathcal{L}_{\mathrm{pred}}^{\mathrm{real}}+\alpha\mathcal{L}_{\mathrm{causal}}^{\mathrm{syn}},$ (8)
where $\alpha$ is a hyperparameter controlling the emphasis on the causal regularizer. Despite its simplicity, we will show in Sec. 5.3 that the sim-to-real causal transfer framework effectively translates the causal knowledge absorbed from simulations to real-world datasets, even without any real-world causal annotations.
@@ 5 Experiments
Throughout our experiments, the multi-agent forecasting task is defined as predicting the future trajectory of the ego agent for 12 time steps, given the history observations of all agents in a scene in the past 8 time steps. Given that our simulator operates deterministically without random variations in agent trajectories, we train our forecasting models to output a single future trajectory for the ego agent, rather than generating multiple potential trajectories in multimodal settings. We evaluate forecasting models on three metrics:
  • Average Displacement Error (ADE): the average Euclidean distance between the predicted output and the ground truth trajectories – a widely used metric to measure the prediction accuracy of a forecasting model.
@@
  • Average Causal Error (ACE): the average difference between the estimated causal effect and the ground truth causal effect – a metric inspired by Roelofs et al. [56] to measure the causal awareness of a learned representation,
@@
Alongside the aggregated ACE, we also measure causal awareness on each category separately, i.e., ACE-NC, ACE-DC, ACE-IC for non-causal, direct causal, and indirect causal agents, respectively. Note that ACE can only be evaluated in the presence of causal annotations $\mathcal{E}_{i}$, e.g., the diagnostic dataset through controlled simulations (Sec. 3.3).
@@
Figure 6: Quantitaive results of our causal regularization method in comparison to the Autobots baseline [22] and the non-causal data augmentation [56]. Models trained by our ranking-based method yield enhanced performance both in terms of both causal effect estimation (left) and prediction accuracy (right). Results are averaged over five random seeds.
(a) ID
(b) OOD-Context
(c) OOD-Density
Figure 7: Qualitative results of our method on in-distribution (ID) and out-of-distribution (OOD) test sets. Models regularized by our ranking-based method show a more robust understanding of agent interactions in high-density and novel scene contexts.
@@ 5.1 Robustness of Recent Representations
Setup. We start our experiments with the evaluation of a collection of recent models on the diagnostic dataset described in Sec. 3. We explicitly examine the errors of causal effect estimation for each agent category. We summarize the results in Tab. 1 and visualize the detailed estimation in Fig. 5.
Takeaway 1: recent representations are already partially robust to non-causal agent removal. As shown in Tab. 1, the values of ACE-NC are generally quite minimal compared to the results on other metrics. Across all evaluated models, the errors made upon non-causal agents are 10x smaller than that for causal agents (ACE-DC/IC). In line with our analysis in Sec. 3, this result provides a counterargument to the robustness issue benchmarked in  Roelofs et al. [56], suggesting the importance of studying the robustness under causal perturbations as opposed to non-causal ones.
@@ 5.2 Effectiveness of Causal Regularization
We next evaluate the efficacy of our causal regularization method on Autobots, the strongest baselines in Tab. 1. As a baseline comparison, we incorporate the data augmentation strategy from [56], which augments training scenarios by selectively removing non-causal agents. This approach assumes that the removal of non-causal agents should not alter the ego agent’s trajectory, thereby using the same ground truth trajectory to reinforce model predictions against changes under non-causal perturbations. We assess the performance in two key areas: in-distribution causal awareness and out-of-distribution generalization.
Figure 8: Results of causal transfer from ORCA simulations to the ETH-UCY dataset. We consider three settings where the model has access to the same simulation data but varying amounts of real-world data. Our ranking-based causal transfer results in lower prediction errors and higher learning efficiency than the other counterparts, i.e., AutoBots baseline [22], the vanilla sim-to-real transfer [38] and the non-causal data augmentation [56]. Details are summarized in Appendices B and A.
@@ 5.2.1 In-distribution Causal Awareness
Takeaway: our method boosts causal awareness thanks to fine-grained annotations. As shown in Fig. 6, both the contrastive and ranking variants of our causal regularization approach result in lower causal errors than the vanilla baseline. In particular, the causal ranking method demonstrates substantial advantages over the other counterparts across all quantities of causal effects, confirming the promise of incorporating fine-grained annotations in learning causally-aware representations.
@@ 5.2.2 Out-of-Distribution Generalization
Setup. We further examine the efficacy of our method by measuring the prediction accuracy on out-of-distribution test sets. We consider two common types of distribution shifts in the multi-agent setting: higher density and unseen context, as detailed in Appendix B. We summarize the results of different methods on the OOD test sets in Fig. 6 and visualize the difference of prediction output in Fig. 7.
Takeaway: causally-aware representations lead to stronger out-of-distribution generalization. As shown in Fig. 6, the prediction error from our causal ranking method is generally lower than that of the other methods on the OOD sets. In fact, the overall patterns between the two parts of  Fig. 6 are highly similar, indicating a substantial correlation between causal awareness and out-of-distribution robustness. This practical benefit is also visually evident in the qualitative visualization in Fig. 7.
@@ 5.3 Effectiveness of Causal Transfer
Setup. Finally, we evaluate the proposed causal transfer approach on the ETH-UCY dataset [36, 52] and NBA rebound dataset [70, 71] paired with our diagnostic dataset. To assess the potential of the transfer method in challenging settings such as low-data regimes, we train Autobots [22] on different fractions of the real-world data. The results are summarized in Fig. 8 and Tab. 2.
Takeaway: causal transfer improves performance in the real world despite domain gaps. As shown in  Fig. 8, our causal transfer approach consistently improves the performance of the learned representation on ETH-UCY, even with a substantial simulator-to-real-world gap. In particular, it enables the model to learn faster, e.g., even with half data in the real-world, our ranking-based method can still outperform other counterparts. In contrast, the vanilla sim-to-real technique [38] which combines synthetic and real data, struggles with effective transfer, especially in low-data regimes. We hypothesize this is due to disparities in non-causal interaction styles, such as speed and curvature, between simulation and reality. We report ORCA simulator and other baselines’ prediction accuracy on ETH-UCY in Appendix A. We further examine the causal transfer approach on NBA rebound, a more large-scale and interactive dataset. Table 2 shows that our ranking-based method outperforms the baseline and demonstrates that the proposed causal transfer can work not only for pedestrians but also for basketball players.
[…]
Baseline 0.562 1.271
Vanilla 0.567 1.287
Augment 0.569 1.292
Contrast (ours) 0.550 1.248
Ranking (ours) 0.544 1.235
Table 2: Results of causal transfer from ORCA simulations to the NBA rebound dataset.We outperform the AutoBots baseline by integrating our causal regularizers, showing effectiveness of our approach. The numbers are in meters.
@@ 6 Conclusions
Summary. In this paper, we presented a thorough analysis and an effective approach for learning causally-aware representations of multi-agent interactions. We revisited the notion of non-causal robustness in the recent benchmark [56] and showed that the main weaknesses of recent representations are not overestimations of the effect of non-causal agents but rather underestimation of the effects of indirect causal agents. To boost causal awareness of the learned representations, we introduced a regularization approach that encapsulates a contrastive and a ranking variant leveraging annotations of different granularity. We showed that our approach enables recent models to learn faster, generalize better, as well as transfer stronger to practical problems, even without real-world annotations.
Discussions. As one of the first steps towards causally-aware representations in the multi-agent context, our work raises two primary questions for further exploration. On the technical front, while our proposed regularization method consistently boosts causal awareness in various settings, supervision alone is likely insufficient to fully solve causal reasoning in complex scenes, as evidenced by the causal errors in Fig. 6. Incorporating structural inductive biases might be a promising direction to address high-order reasoning challenges in the presence of indirect causal effects.
7 Acknowledgement
The authors would like to thank Yang Gao and Kaouther Messaoud for their valuable feedback. This work was supported by Honda R&D Co., Ltd and Hasler Foundation under the responsible AI program.